OpenStudy (anonymous):

Let f(x) = x ^3. The equation of the tangent line to f(x) at x = 5 is y= . Using this, we find our approximation for 4.8 ^3 is .

OpenStudy (anonymous):

I'm so confused

OpenStudy (sosa4954):

first inver opperation

OpenStudy (anonymous):

what do you mean sir?

OpenStudy (sosa4954):

im a girl

OpenStudy (perl):

The equation of the tangent line to f(x) at x = a is L(x) = f(a) + f '(a) * ( x - a )

OpenStudy (sosa4954):

there you go

OpenStudy (perl):

f(x) = x^3 f ' (x) = 3x^2

OpenStudy (perl):

f(5) = 5^3 = 125 f ' (5) = 3*5^2 = 75

OpenStudy (perl):

L(x) = f(a) + f '(a) * ( x - a ) L(x) = 125 + 75 ( x - 5 )

OpenStudy (perl):

now plug in x = 4.8 L(4.8) = 125 + 75 ( 4.8 - 5) = 110 this is very close to the exact answer of 4.8^3 = 110.592

OpenStudy (anonymous):

thank you so much @perl